OpenStudy (anonymous):

Find a formula for F^(n)(x) if f(x)=e^(2X+1) f^(n)(x)=?

OpenStudy (raden):

u mean the generalize derivative of f(x) right ?

OpenStudy (anonymous):

i think so.

OpenStudy (anonymous):

i think i have to take derivatives like 3 or 4 times and come up with a formula

OpenStudy (raden):

for this case, we can use the chail rule f(x)=e^(2X+1) ---> f ' (x) = 1/2 * e^(2X+1) it is the first derivative

OpenStudy (raden):

*chain

OpenStudy (raden):

f '(x)=1/2 * e^(2X+1) --->f '' = 1/2*1/2*e^(2X+1) = (1/2)^2*e^(2x+1) it is the 2nd derivative

OpenStudy (anonymous):

i thought it was this first derivative 2e^(2x+1) second 4e^(2x+1) third 8e^(2x+1)

OpenStudy (raden):

opppss... i done it by integration, haha my bad sorry, u are right first derivative 2e^(2x+1) second 4e^(2x+1) = (2)^2e^(2x+1) third 8e^(2x+1) = (2)^3e^(2x+1) .... and so on we'll get the generalize for derivative of f(x) : f^(n)(x) = (2)^n * e^(2x+1)

OpenStudy (anonymous):

perfect!!!

OpenStudy (anonymous):

how did you do the generalizing?

OpenStudy (raden):

thanks for medal, and for this time i have to give u medal too :)

OpenStudy (anonymous):

haha :) tnx

OpenStudy (anonymous):

i see i see it now, i was trying 2n, but it was 2^n

OpenStudy (raden):

look, the rithme the numbers 2,4,8,16,32,64,... so on it can be 2, 2^2, 2^3, 2^4, 2^5, 2^6, ..., till 2^n

OpenStudy (anonymous):

got it!

OpenStudy (raden):

yeah, u are right